OpenStudy (anonymous):

Find the midpoint of the segment with the given endpoints. M(–2, 1) and O(–3, 2) A. (–2.5, 1.5) B. (–0.5, –0.5) C. (0.5, –0.5) D. (–3, –5)

OpenStudy (hoblos):

x= (-2-3)/2 = -5/2 =-2.5 y= (1+2)/2 = 3/2 = 1.5

OpenStudy (anonymous):

Average the x values, average the y values.

OpenStudy (hoblos):

A

OpenStudy (anonymous):

midpoint x coordinate=(x1+x2)/2 y coordinate =(y2+y1)/2 now calculate